You are here: Foswiki>Commissioning Web>Com20090205 (05 Feb 2009, AndrewRakich)Edit Attach

BG February Com Night 5, February 5 2009

Observers: A. Rakich J. Hill
Telescope Operator: D. Gonzales Huerta
SW Support: Name
Telescope Support: Name

Summary:

More field data unnecessary. Checked the patch to presets in PSF, worked well. Got some more "fast" WFS data in good seeing. Running with new MCSPU with no problems.

Preparations:

Tonight will be spent tuning up the new MCSPU software. In the end, the goal is to hand over a version of the code that we consider to be stable. We may deploy the new axis controller or we may revert back to the old version depending on how stable the new software appears to be.

TCS Build 20 was installed this afternoon. Build 20 has the new preset code for M2 Offload Compensation, and has a LUCIFER compatible IIF.

10:25 send preset to BS9145 in acquire mode, adjust IE=-39 CA=-7 resend preset in active mode, collimating

10:43 collimating on BS9142 converged to 100+ nm in the PSF M1 GUI we see that Z22 is being sent and updated, but in the GCS GUI onlt Z4-11 is being reported. This was not happening last week.

10:55 doing experiment to check that the preset collimation error is now fixed; collimating in split mode with M2 doing Z11 and Z22 focus components , collimated for 10 minutes on BS9142 @ 59 deg, now going to BS9148 @85 degrees, note that the star is still well collimated so that worked well. John gives bonus points to Torsten for acquiring a star at the edge of the field, but subtracts half after noticing that the acuisition image is still not fully displayed to the edge...

11:01 guiding suddenly hung, stopped receiving images from AZCAM, killed GCS and resent preset which got things going again.

11:07 starting fast WFS measurements, images in 090205w_fast,

(note the ~1200 nm drifits in focus for each set of 12 5 second measurements probably due to the mirror ventillation being switched off).

seeing 0.5"-0.7" blowing up towards the end.

collimating 2-12 12X5sec exposures 13-24 collimating 25-30 12X5sec exposures 31-42

interrupt to quickly check offset functionality; offset worked well.

collimating 43-51 12X5sec exposures 52-63 collimating 64-71 aborting because of seeing blow up.

11:56 Changing the guide rate

Previous guide rate 0.2 effective guide rate 0.17 now changing to 0.5; effective guide rate reported as 0.15 - 0.23 set guide rate to 1, effective rate reported 0.15-0.23 12:01 noticed that just as seeing blew up, the image got a lot more spread out, the exposure time dropped !!!. Something's funny there.

12:05 Roberto calls, and John takes over to do single eye LBCR picture of comet Lulin

Details (Times in MST):

Open

HH:MM Open Dome, T=? degC. D=? degC, Wind (LBTO) ? m/sec.

Close

HH:MM Close dome T=?degC, D=?degC, wind (LBTO) ? m/sec.

-- JoarBrynnel - 15 Jan 2009
Topic revision: r4 - 05 Feb 2009, AndrewRakich
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback